A student eat a dinner containing 8.0×10^6J of energy and wishes to lift an object in a gym of 60kg .How mang times will he lift this object to expand this much energy if he lift this object in a distance of 2.0m


1
Expert's answer
2022-03-22T15:34:59-0400

The energy change is equal to work done

"E=Nmgh"

Hence, the number of lifting of the weight

"N=\frac{E}{mgh}=\frac{8.0\times 10^6}{60\times 9.8\times 2.0}=6803"